OpenStudy (anonymous):

the measure of one angle of a triangle is 40. what are the measures of the other two angles in the ration of 3:4

OpenStudy (anonymous):

60 and 80

OpenStudy (anonymous):

180 - 40 = 140 140/7 = 20 where 7 --> (4+3) from the ratio Then the 4*20 =80 and the 3 *20 = 60

OpenStudy (anonymous):

or think of it as 40 + 3x + 4x = 180